OpenStudy (anonymous):

What is the sum of the arithmetic sequence 5, 11, 17 …, if there are 24 terms?

OpenStudy (cwrw238):

common difference d = 11-5 = 6 sum of n terms = (n/2)[2a + (n-1)d] plug in n = 24, a = 5 and d = 6 to get your answer

OpenStudy (shubhamsrg):

i'll tell you a completely different method first calculate the 24th term.. it will be 5+(23)(6) = 143 now keep adding 1st and last terms,,move from left to right (from beginning) and from right to left ( from end) i.e. 5+143 = 148 11 + 137 =148 17 + 131 =148 . . . 12 times add all eqns..so your ans is 148*12 !!

OpenStudy (shubhamsrg):

you can also see it as sum of n terms = n(first term + last term)/2 or you may use the formulla which cwrw told..its the same thing..

OpenStudy (anonymous):

thank youuu!